what is 3 to the tenth power divided by 3 to the second power?
Answer by stanbon(75887)

what is 3 to the tenth power divided by 3 to the second power?
-----
You have 3 as a factor 10 times in the numerator
and 3 as a factor 2 times in the denominator.
----
Cancel 2 of the factors in the numerator and
2 in the denominator. That leaves you with
8 3's in the numerator.
So the answer is 10^8.
---
The general rule is 3^10/3^2 = 3^(10-2) = 3^8
================
Cheers,
Stan H.